OpenStudy (anonymous):

What polynomial has roots of –4, 1, and 6 ?

jimthompson5910 (jim_thompson5910):

x = -4, x = 1 or x = 6 x+4 = 0 or x-1 = 0 or x-6 = 0 (x+4)(x-1)(x-6)=0 (x+4)(x^2-7x+6)=0 x(x^2-7x+6)+4(x^2-7x+6)=0 x^3-7x^2+6x+4x^2-28x+24=0 x^3-3x^2-22x+24=0 So the polynomial x^3-3x^2-22x+24 has the roots –4, 1, and 6

OpenStudy (anonymous):

medal for you

OpenStudy (anonymous):

woot i got the same thing XD....why is always the easy ones i mees up on